# Right Splitting, Galois Correspondence, Galois Representations and Inverse Galois Problem

### v1

@@ 3.3. Direct sum / tensor product of representations and Galois groups
Now $H\subset G^{\prime}\subset G$. Since $G=\bigsqcup_{1\leq i\leq n}Hs_{i}$, we have $G^{\prime}=\bigsqcup_{i\in I}(G^{\prime}\cap Hs_{i})$ where $I=\{i\in\{1,2,...,n\}\ |\ G^{\prime}\cap Hs_{i}\neq\emptyset\}$. Since $H\subset G^{\prime}$, we have $(G^{\prime}\cap Hs_{i})\neq\emptyset\iff s_{i}\in G^{\prime}$. Hence $I=\{i\in\{1,2,...,n\}\ |\ s_{i}\in G^{\prime}\}$. Now for $i\in I$, it is easy to see that, $(G^{\prime}\cap Hs_{i})=(G^{\prime}\cap H)(s_{i})=Hs_{i}$. Hence $G^{\prime}=\bigsqcup_{i\in I}Hs_{i}$ and $G^{\prime}/H=\{s_{i}H\}_{i\in I}$. Now the following argument completes the proof.
$\begin{split}&ker(\Phi)=G^{\prime}/H\ \text{under}\ \Omega\\ \iff&\{s_{i}H\}_{i\in J}=\{s_{i}H\}_{i\in I},\text{that is}\ I=J\\ \iff&s_{i}\in G^{\prime}\ \text{iff}\ ker(\pi)\cap Hs_{i}\neq\emptyset.\\ \iff&s_{i}\in G^{\prime}\ \text{ iff there is}h_{i}\in H\ \text{such that}\pi(h_{i}s_{i})=I.\\ \iff&s_{i}\in G^{\prime}\ \text{ iff}\ \pi(s_{i})\in\pi(H)\\ \iff&g\in G^{\prime}\ \text{ iff }\ \pi(g)\in\pi(H)\\ \iff&G^{\prime}\ \text{ is largest subgroup of}G\ \text{such that}\ \pi(G^{\prime})=\pi(H).\end{split}$
@@
$1\rightarrow ker(\Phi)\xhookrightarrow{}L\rightarrow\pi(G)\rightarrow 1$
is right split. $L\cong\pi(G)\ltimes G/H$ for some semidirect product group law.
@@
Since $\pi(G)=\pi(H)$, $ker(\Phi)=G/H$ under $\Omega$.
Consider $\iota:\pi(G)\rightarrow L$ given by $\iota(x)=(\pi(h),\rho(h))$ where $x=\pi(h)$ for some $h\in H$.
If $x=\pi(h)=\pi(h^{\prime})$ for some $h,h^{\prime}\in H$ then $\pi(hh^{\prime-1})=I$. Hence by matrix calculation above, $\rho(hh^{\prime-1})=I$. Hence $\rho(h)=\rho(h^{\prime})$. Thus $\iota$ is well defined.
Now, for $x,y\in\pi(G)$, let $h,h^{\prime}\in H$ such that $x=\pi(h),y=\pi(h^{\prime})$. Then
$\iota(x)\iota(y)=(\pi(h),\rho(h))(\pi(h^{\prime}),\rho(h^{\prime}))=(\pi(hh^{\prime}),\rho(hh^{\prime}))=\iota(\pi(hh^{\prime}))=\iota(\pi(h)\pi(h^{\prime}))=\iota(xy).$
Hence $\iota$ is a homomorphism. Also since, $\Phi\circ\iota=id_{\tilde{\pi}(G)}$, the exact sequence splits.
@@
Let $G/H$ be cyclic $<sH>$. Then if $ker(\pi)\cap Hs\neq\emptyset$ then the above exact sequence is right split.
@@
Since $ker(\pi)\cap Hs\neq\emptyset$, there is an $h_{1}\in H$ such that $\pi(h_{1}s)=1$. Since $H$ is normal in $G$, $(h_{1}s)^{i}=h_{i}s^{i}$ for some $h_{i}\in H$. Hence $\pi(h_{i}s^{i})=\pi(h_{1}s)^{i}=1$. Hence for all $i$, $ker(\pi)\cap Hs^{i}\neq\emptyset$. Thus $ker(\Phi)\cong G/H$.
@@
Let $\{\rho(s_{i})\}_{i}$ form a multiplicatively closed subset of $\rho(G)$ and let $(n,|\pi(G)|)=1$. Then above exact sequence is right split.
@@
Since $\tilde{\pi}(H)=\tilde{\pi}(G)$ as in Thm 3.12, we have from above that ${\rm PSL}_{2}(\mathbb{F}_{p})\ltimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law as in 3.19). Since $Aut(\mathbb{Z}/2\mathbb{Z})$ is trivial, we have that semidirect product is indeed direct. ∎
We can generalize Lemma 3.16, Corollary 3.17.1, Lemma 3.18 and Thm 3.19. Consider a closed subgroup $H^{\prime}$ of $G$. We have $H^{\prime}/(H^{\prime}\cap H)\xhookrightarrow{}G/H$.
@@
    $1\rightarrow ker(\xi)\rightarrow N\rightarrow\pi(H^{\prime})\rightarrow 1$
    is right split. $N\cong\pi(H^{\prime})\ltimes H^{\prime}/(H^{\prime}\cap H)$ for some semidirect product group law.
@@
    Let $H\subset H^{\prime}$. Suppose $\pi(G)=\pi(H)$. Then the above exact sequence is right split. $N\cong\pi(H^{\prime})\ltimes H^{\prime}/H$ for some semidirect product group law.

### v2

@@ 3.3. Direct sum / tensor product of representations and Galois groups
Now $H\subset G^{\prime}\subset G$. We have $G^{\prime}=\bigsqcup_{i\in I}(G^{\prime}\cap Hs_{i})$ where $I=\{i\in\{1,2,...,n\}\ |\ G^{\prime}\cap Hs_{i}\neq\emptyset\}$. Also $(G^{\prime}\cap Hs_{i})\neq\emptyset\iff s_{i}\in G^{\prime}$. Hence $I=\{i\in\{1,2,...,n\}\ |\ s_{i}\in G^{\prime}\}$. For $i\in I$, $(G^{\prime}\cap Hs_{i})=Hs_{i}$. Hence $G^{\prime}=\bigsqcup_{i\in I}Hs_{i}$ and $G^{\prime}/H=\{s_{i}H\}_{i\in I}$. Now the following argument completes the proof.
[…]
@@
(3.3) $1\rightarrow ker(\Phi)\xhookrightarrow{}L\rightarrow\pi(G)\rightarrow 1$
is right split and $L\cong\pi(G)\ltimes G/H$ for some semidirect product group law.
@@
Since $\pi(G)=\pi(H)$, $ker(\Phi)=G/H$ under $\Omega$. Consider $\iota:\pi(G)\rightarrow L$ given by $\iota(x)=(\pi(h),\rho(h))$ where $x=\pi(h)$ for some $h\in H$. If $x=\pi(h)=\pi(h^{\prime})$ for some $h,h^{\prime}\in H$ then $\pi(hh^{\prime-1})=I$. Hence by matrix calculation above, $\rho(hh^{\prime-1})=I$. Hence $\rho(h)=\rho(h^{\prime})$. Thus $\iota$ is well defined. Clearly, $\iota$ is a homomorphism and $\Phi\circ\iota=id_{\tilde{\pi}(G)}$, therefore the exact sequence splits.∎
@@
Let $G/H$ be cyclic <sH><sH>. If $ker(\pi)\cap Hs\neq\emptyset$ then the exact sequence 3.3 is right split.
@@
Since $ker(\pi)\cap Hs\neq\emptyset$, there is an $h_{1}\in H$ such that $\pi(h_{1}s)=1$. Since $H$ is normal in $G$, $(h_{1}s)^{i}=h_{i}s^{i}$ for some $h_{i}\in H$. Hence $\pi(h_{i}s^{i})=\pi(h_{1}s)^{i}=1$. Hence for all $i$, $ker(\pi)\cap Hs^{i}\neq\emptyset$. Thus $ker(\Phi)\cong G/H$.∎
@@
Let $\{\rho(s_{i})\}_{i}$ form a multiplicatively closed subset of $\rho(G)$ and let $(n,|\pi(G)|)=1$. Then the exact sequence 3.3 is right split.
@@
Since $\tilde{\pi}(H)=\tilde{\pi}(G)$ as in Theorem 3.12, we have from above that ${\rm PSL}_{2}(\mathbb{F}_{p})\ltimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law as in 3.19). Since $Aut(\mathbb{Z}/2\mathbb{Z})$ is trivial, we have that semidirect product is indeed direct. ∎
We can generalize Lemma 3.16, Corollary 3.17.1, Lemma 3.18 and Theorem 3.19. Consider a closed subgroup $H^{\prime}$ of $G$. We have $H^{\prime}/(H^{\prime}\cap H)\xhookrightarrow{}G/H$.
@@
    (3.4) $1\rightarrow ker(\xi)\rightarrow N\rightarrow\pi(H^{\prime})\rightarrow 1$
    is right split and $N\cong\pi(H^{\prime})\ltimes H^{\prime}/(H^{\prime}\cap H)$ for some semidirect product group law.
@@
    Let $H\subset H^{\prime}$. Suppose $\pi(G)=\pi(H)$. Then the exact sequence 3.4 is right split and $N\cong\pi(H^{\prime})\ltimes H^{\prime}/H$ for some semidirect product group law.